top of page

Depot Report - Saturday 17th February 2024


55009

No work on the loco this week. It is due to be tested on the Barrow Hill demonstration line in the coming days.




55015

Work continued on sorting a sticking door at No 2 end, as well as other miscellaneous items. A new collector drum drain tap was fitted, as well as putting new pipework in for the screen wash. Volunteers were sanding down and repainting 15's bodywork while continuing with the painting of the steam heat boiler out of Tulyar.


L


55019

Work started of by wiping down the paintwork. Volunteers were in the pit securing the connections and cables in readiness for the move this week. Other volunteers have replaced the traction motor bellows with a new one. Finally, work progressed with sorting out the remaining electrical issues on Royal Highland Fusilier.



Thanks to volunteers Mike, Geoff, Peter, Richard, Bill. Steve, Simon, Amanda, Alex, Chris, Daniel & Nigel.


Photos: Howard Jones

179 views0 comments

Recent Posts

See All

Comments


bottom of page